OpenStudy (brianna23x):

find the distance between these two points: (-3, -4) and (0,0)

Mehek (mehek14):

\(\large{\bf{\sqrt{(0--4)^2+(0--3)^2}}}\)

OpenStudy (brianna23x):

is it -5, 22 squared, or 5?

Mehek (mehek14):

Well what's 0 - (-4) ?

OpenStudy (brianna23x):

-4?

Mehek (mehek14):

no 0 - (-4) = 0 + 4 = ?

OpenStudy (brianna23x):

4 = 4 = ?

OpenStudy (anthonyym):

You can draw a picture. The distance is the hypotenuse. Find the hypotenuse using Pythagorean theorem.

OpenStudy (mathstudent55):